# Measurement of microwave photon correlations at millikelvin with a thermal detector

### v1

@@ Abstract
Figure 4 summarizes the correlation measurement result as well as the raw data of $\langle n\rangle$ and $(\Delta n)^{2}$. At a fixed power of the coherent field, $P_{\rm h}$, the correlation function, $g^{(2)}(0)$, increases with the radiation temperature, $T$, from $1$ to a value slightly below $2$. It reveals the loss of coherence as the thermal proportion of the field increases. On the other hand, $g^{(2)}(0)$ converges to $1$ with increasing $P_{\rm h}$. This is consistent with our expectation since an ideal coherent field leads to $g^{(2)}(0)=1$. The measured value of $g^{(2)}(0)$ is larger than the theoretical expectation by a few percentage where the coherent photon number is below $1\,{\rm photon/(s\times Hz)}$. We attribute this loss of coherence to the approximately $80\,{\rm dB}$ attenuation in the input line, where the blackbody radiation at different thermalization stages necessarily impinges on the input signal. Overall, the observed $g^{(2)}(0)$ shows a quantitative agreement with the theoretical expectation. It is directly obtained from the resonance frequency and linewidth of the averaged scattering response without amplification and the corresponding noise. This is the key advantage of using cryogenic detectors for microwave photon correlation measurement compared with the combination of quantum limited amplifiers and room-temperature linear detectors [31, 32].
@@
The bolometric measurement of the second-order photon correlation function demonstrates the ability of using cryogenic detectors to characterize the quantum statistical properties of microwave radiation. This method performs the projective quantum measurement directly at the millikelvin stage, and therefore only requires a data bus from the cryogenic to the room temperature stages to transfer the classical measurement outcome. Multiple nanobolometers can be multiplexed with the same bus. The combination of long and short SNS junctions allows the RF operation of the nanobolometer, which plays a key role in reading out the second-order photon number moments. This configuration also pushes the bandwidth of cryogenic detectors to the microwave regime as compared to other designs such as transition edge sensor (TES) [33], kinetic inductance detector (KID) [34], and superconducting nanowire single photon detector (SNSPD) [35]. By combining this technique with correlation spectroscopy [26], it is possible to measure even higher-order correlation functions of microwave photons with a single nanobolometer. An alternative upgrade for higher-order correlation measurements is to combine the nanobolometers with microwave beam splitters [36] in a similar way of the well established optical experiment setup. In addition to the advantage of a small physical footprint, integration of the microwave photon generators and detectors at the cryogenic temperature enables an isolated cryogenic environment for quantum information processors. It opens up new possibilities of controlling and reading out the quantum information carried by microwave photons in a more efficient and timely fashion [37, 38].
@@ Methods
The nanobolometer used in this study consists of an $7.5\,{\rm mm}$-long Nb coplanar waveguide filter, an $200\,{\rm nm}$-wide $1\,{\rm\mu m}$-long Au3Pd absorber, and an $7$-island Al-Au3Pd-Al junction thermometer (see Extended Data Fig. 1) [22, 23]. The superconducting islands are $300\,{\rm nm}$ wide and $2.4\,{\rm\mu m}$ long, and they are evenly spaced by a $300\,{\rm nm}$ gap. The designed thicknesses of the Si substrate and the SiO2, Nb, Au3Pd, and Al layers are $675\,{\rm\mu m}$, $300\,{\rm nm}$, $200\,{\rm nm}$, $30\,{\rm nm}$, and $100\,{\rm nm}$, respectively.
@@
The cryogenic and room-temperature setup for the experiment is shown in Extended Data Fig. 2. We use a directional coupler at the thermometer input for reflection-type measurement, where the input signal is attenuated by approximately $20\,{\rm dB}$ but almost all the signal reflected from the sample is routed to the output port. The signal is subsequently amplified and down converted to an intermediate frequency of $62.5\,{\rm MHz}$ for measurement. Here, the filters before and after the mixers are used to remove the red sideband. The sampling rate is $250\,{\rm Msps}$. We obtain one data point of the IQ quadratures in each $16\,{\rm ns}$ period, which is then digitally filtered by a $500\,{\rm kHz}$ low-pass finite-impulse-response filter before averaging. The repetition rate of the measurement is $1.25\,{\rm kHz}$. In each repetition, we obtain a $32\,{\rm\mu s}$-long trace of the IQ quadratures, which is subsequently averaged by $2\times 10^{4}$ times.
@@
The reflection coefficient of the thermometer is described as $S_{11}=1-\textrm{e}^{\textrm{i}\phi}\gamma_{\rm c}/\left[(\gamma/2)+\textrm{i}\Delta\right]$[39], where $\gamma_{\rm c}$ and $\gamma$ are the external and the total energy decay rates, and $\Delta=2\pi(f_{\rm r}-f_{\rm p})$ is the detuning between the resonance and the probe frequencies. The parameter $\phi$ describes the asymmetry of the resonance. We consider a Gaussian distribution of the resonance frequency, $f_{\rm r}\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$, which is caused by the statistical properties of the input photons. The averaged thermometer response is thus given by
$\langle S_{11}\rangle=1-\frac{\textrm{e}^{\textrm{i}\phi}\gamma_{\rm c}}{2\sqrt{2\pi}\sigma}{\rm erfcx}\left(\frac{(\gamma/2)+\textrm{i}\Delta^{\prime}}{2\sqrt{2}\pi\sigma}\right).$ (1)
Here, $\Delta^{\prime}=2\pi(\mu-f_{\rm p})$ and ${\rm erfcx}(\cdot)$ is the scaled complementary error function. The magnitude of $\langle S_{11}\rangle$ is equivalent to the Voigt profile in laser spectroscopy [26].
0.8 Fitting procedure.
We model the thermometer as a parallel RLC circuit, described by resistance $R$, inductance $L$, and capacitance $C$, which is capacitively coupled to the feedline via a small capacitor, $C_{\rm g}$. The characteristic impedance of the feedline is assumed to be $Z_{0}=50\,{\Omega}$. Assuming that $2\pi f_{\rm r}ZC_{\rm g}\ll 1$ with $Z=\sqrt{L/C}$, we expand the input impedance in a Taylor series around the resonance frequency, $f_{\rm r}$, and obtain the input impedance as $Z_{\rm in}=R^{\prime}-i2L^{\prime}\Delta f$[1]. Here, $R^{\prime}=1/(8\pi ZC_{\rm g}^{2}Q_{\rm i}f_{\rm r}^{2})$, $L^{\prime}=1/(8\pi ZC_{\rm g}^{2}f_{\rm r}^{3})$, and $Q_{\rm i}$ is the internal quality factor of the RLC circuit. The external quality factor is $Q_{\rm e}\equiv Q_{\rm i}R^{\prime}/Z_{0}$. Thus, we obtain the internal and external damping rates as $\gamma_{\rm i}=2\pi f_{r}/Q_{\rm i}$ and $\gamma_{\rm c}=4ZZ_{0}C_{\rm g}^{2}f_{\rm r}$[39]. Assuming that $Q_{\rm i}$ is a constant and $f_{\rm r}$ varies in a relatively small range compared to itself, we treat $\gamma_{\rm i}$ to be independent of $f_{\rm r}$ while $\gamma_{\rm c}$ as a fitting parameter in the ideal reflection response, $\langle S_{11}\rangle$.
We phenomenally describe the output path by the transfer function $H(f_{\rm p})=s_{\rm b}+e^{i\phi_{\rm b}}\gamma_{\rm b,c}/[(\gamma_{\rm b}/2)+i\Delta_{\rm b}]$, in order to extract the actual signal from the $80\,{\rm MHz}$-period Lorentzian patterns in the output signal (see Extended Data Fig. 3a). It possibly arises from a large mismatch between the sample and the HEMT amplifier (see Extended Data Fig. 2). Here, $s_{\rm b}$ is a scaling factor, $\Delta_{\rm b}=2\pi(f_{\rm b}-f_{\rm p})$ is the detuning between the resonance and probe frequencies, $\gamma_{\rm b,c}$ and $\gamma_{\rm b}$ are free parameters of a Lorentzian function, and $\phi_{\rm b}$ is the asymmetry factor of the lineshape. Considering also the cable delay, $\tau$, and phase offset, $\varphi$, the output signal is $S_{11}(f_{\rm p})=e^{i(f_{p}\tau+\varphi)}H(f_{\rm p})\langle S_{11}(f_{\rm p})\rangle$. In total, there are $10$ free parameters to be determined before extracting $\mu$ and $\sigma$. We first fit all the parameters at the base temperature and fix $6$ of them throughout the experiment (see Extended Data Fig. 3b-c). The only fitting parameters in each individual measurement are $\gamma_{\rm c}$, $\phi$, $f_{\rm b}$, and $\phi_{\rm b}$, as shown in Extended Data Fig. 3d. The fitted results are almost identical for thermal and coherent inputs with maximally few percentage of difference at the highest temperatures ($T>1.7\,{\rm K}$).
@@ Acknowledgments
Extended Data Fig. 3: Fitting procedure for second-order correlation. a Magnitude of the reflection coefficients, where we observe spurious resonances with a $80\,{\rm MHz}$ spacing possibly due to a large impedance mismatch in the output path. b The magnitude and phase of the reflection coefficients (blue circles) together with the fitting results (red curves) in a smaller frequency range, shown as the gray area of (a). c The fitted ideal reflection coefficient, $\langle S_{11}(f_{\rm p})\rangle$, and the transfer function, $H(f_{\rm p})$. The legend shows the fixed parameters of the model throughout the experiment (see Methods). d The fitted parameters of the model for coherent (triangle) and thermal (circle) input fields. The red triangular markers are the specific values for (b) and (c).

### v2

@@ Results
Figure 4 summarizes the correlation measurement result as well as the raw data of $\langle\hat{n}\rangle$ and $(\Delta n)^{2}$. At a fixed power of the coherent field, $P_{\rm h}$, the correlation function, $g^{(2)}(0)$, increases with the radiation temperature, $T$, from $1$ to a value slightly below $2$. It reveals the loss of coherence as the thermal proportion of the field increases. On the other hand, $g^{(2)}(0)$ converges to $1$ with increasing $P_{\rm h}$. This is consistent with our expectation since an ideal coherent field leads to $g^{(2)}(0)=1$. The measured value of $g^{(2)}(0)$ is larger than the theoretical expectation by a few percent where the coherent photon number is below $1\,{\rm photons/(s\times Hz)}$. We attribute this loss of coherence to the $88\,{\rm dB}$ attenuation in the input line, where the blackbody radiation at different thermalization stages necessarily impinges on the input signal. Overall, the observed $g^{(2)}(0)$ shows a quantitative agreement with the theoretical expectation. It is directly obtained from the resonance frequency and linewidth of the averaged scattering response without amplification and the corresponding noise. This is the key advantage of using cryogenic detectors for microwave photon correlation measurement compared with the combination of quantum limited amplifiers and room-temperature linear detectors [33, 34].
Discussion
@@
The bolometric measurement of the second-order photon correlation function demonstrates the ability of using cryogenic detectors to characterize the quantum statistical properties of microwave radiation. This method implements projective quantum measurements directly at the millikelvin stage, and therefore only requires a data bus from the cryogenic to the room temperature stages to transfer the classical measurement outcome. Multiple nanobolometers can be multiplexed with the same bus [35]. The combination of long and short SNS junctions allows the RF operation of the nanobolometer, which plays a key role in reading out the second-order photon number moments. This configuration also pushes the bandwidth of cryogenic detectors to the convenient sub-$10\,{\rm GHz}$ regime as compared to other designs such as transition edge sensor (TES) [36], kinetic inductance detector (KID) [37], and superconducting nanowire single photon detector (SNSPD) [38]. By combining this technique with correlation spectroscopy [26], it is possible to measure even higher-order correlation functions of microwave photons with a single nanobolometer. An alternative upgrade for higher-order correlation measurements is to combine the nanobolometers with microwave beam splitters [39] in a similar way of the well established optical experiment setup.
In the future, the capability of microwave correlation measurements may advance the current understandings of the cosmic microwave background (CMB) [40] and resolve a possible distortion of the CMB frequency spectrum from a pure blackbody radiator. In addition, the small physical footprint and ultralow power consumption of the nanobolometer [23, 24] make it suitable for millikelvin integration. Integrating high-performance microwave photon generators [41] and detectors [42] at millikelvin enables an isolated cryogenic environment for quantum information processors. It opens up new possibilities of controlling and reading out the quantum information carried by microwave photons with less noise, latency, and power consumption.
@@ Methods
The nanobolometer used in this study consists of an $7.5\,{\rm mm}$-long Nb coplanar waveguide filter, an $150\,{\rm nm}$-wide $1\,{\rm\mu m}$-long Au3Pd absorber, and an $7$-island Al-Au3Pd-Al junction thermometer (see Extended Data Fig. 1) [22, 23]. The superconducting islands are $300\,{\rm nm}$ wide and $2.4\,{\rm\mu m}$ long, and they are evenly spaced by a $300\,{\rm nm}$ gap. The designed thicknesses of the Si substrate and the SiO2, Nb, Au3Pd, and Al layers are $675\,{\rm\mu m}$, $300\,{\rm nm}$, $200\,{\rm nm}$, $30\,{\rm nm}$, and $100\,{\rm nm}$, respectively.
@@
The cryogenic and room-temperature setup for the experiment is shown in Extended Data Fig. 2. We use a directional coupler at the thermometer input for reflection-type measurement, where the input signal is attenuated by approximately $20\,{\rm dB}$ but almost all the signal reflected from the sample is routed to the output port. The signal is subsequently amplified and down converted to an intermediate frequency of $62.5\,{\rm MHz}$ for measurement. Here, the filters before and after the mixers are used to remove the red sideband. The sampling rate is $250\,{\rm MS/s}$. We obtain one data point of the IQ quadratures in each $16\,{\rm ns}$ period, which is then digitally filtered by a $500\,{\rm kHz}$ low-pass finite-impulse-response filter before averaging. The repetition rate of the measurement is $1.25\,{\rm kHz}$. In each repetition, we obtain a $32\,{\rm\mu s}$-long trace of the IQ quadratures, which is subsequently averaged by $2\times 10^{4}$ times.
@@
The instantaneous reflection coefficient of the thermometer is described as $S_{11}=1-\textrm{e}^{\textrm{i}\phi}\gamma_{\rm c}/\left[(\gamma/2)+\textrm{i}\Delta\right]$[43], where $\gamma_{\rm c}$ and $\gamma$ are the external and the total energy decay rates, and $\Delta=2\pi(f_{\rm r}-f_{\rm p})$ is the detuning between the resonance and the probe frequencies. The parameter $\phi$ describes the asymmetry of the resonance. We consider a Gaussian distribution of the resonance frequency, $f_{\rm r}\sim\mathcal{N}\left(\mu,\sigma^{2}\right)$, which is caused by the statistical properties of the input photons (see Supplementary Note 2). The averaged thermometer response is thus given by
$S_{11}^{\rm ave}=1-\frac{\textrm{e}^{\textrm{i}\phi}\gamma_{\rm c}}{2\sqrt{2\pi}\sigma}{\rm erfcx}\left(\frac{(\gamma/2)+\textrm{i}\Delta^{\prime}}{2\sqrt{2}\pi\sigma}\right).$ (1)
Here, $\Delta^{\prime}=2\pi(\mu-f_{\rm p})$ and ${\rm erfcx}(\cdot)$ is the scaled complementary error function (see Supplementary Note 3). The magnitude of $S_{11}^{\rm ave}$ is equivalent to the Voigt profile in laser spectroscopy [26]. Considering the contribution of the external circuitry, the real measured reflection coefficient is denoted as $S_{11}^{\rm raw}$ (see Supplementary Note 4).